### v1

@@ 1 Introduction
Traditional neural networks (NNs) have shown promise in analyzing cardiac movement from imaging data [3] through the use of designated architectures [13][15][19][25], optimization schemes [2], priors and constraints [26][5]. However, in addition to the scarcity of annotated high-resolution imagery, they also often suffer from the locality problem, since they struggle to predict tangential movements accurately (as depicted in figure 1(a)). While they excel at minimizing photometric energy to capture radial movements of the myocardium, their inherent locality limits their ability to comprehensively model subtle tangential movements. This limitation is referred to as the aperture problem, as illustrated in figure 1(b).
To address the challenge of capturing both radial and tangential components of myocardium motion, we propose a holistic approach that integrates DL techniques with spectral methods, and specifically FMs [16]. By considering the LV surface as a manifold, we leverage FMs’ global perception for understanding myocardium motion. Our comprehensive approach as detailed in this paper includes a 3D DL architecture and the incorporation of spectral methods. We leverage 2D constraints derived from spectral correspondence methods to significantly enhance the accuracy of surface mapping. These constraints provide a more robust foundation compared to traditional segmentation priors by precisely specifying the pixel correspondences between timesteps.
@@ 3 The Aperture Problem
Global approaches like Horn and Schunk [6],and Convolutional Neural Networks (CNNs) have been suggested but remain curtailed by the aperture problem. Despite CNNs’ broader perspective, their local context makes them prone to this issue. In cardiac CT imaging, this limitation is a major impediment because CNNs may fail to accurately estimate tangential motion on the dynamic LV myocardium surface due to their localized receptive fields.
@@ 4 Method
Figure 2: CardioSpectrum Architecture: The NN analyzes systole and diastole 3D images while incorporating 2D constraints from ZoomOut. These constraints are generated by converting segmentations from both the systole and the diastole into meshes for ZoomOut. The outcome is a 3D optical flow.
@@ 4.1 Global Constraints from Functional Maps
To extract constraints over the LV surface, we leverage the ZoomOut method [12], which iteratively assesses the FM matrix, by transitioning between spatial and spectral domains. This process establishes point-to-point correspondences between the $systole$ and the $diastole$ triangle mesh representations. The meshes, derived from smoothed LV segmentations, serve as a foundation for generating constraints. To map surface constraints to a 3D grid, we calculate an array of 3D displacement, denoted as $d$, between corresponding mesh vertices, as inferred from the ZoomOut results.
By utilizing rounded vertex locations as indices, we fill a 3D grid-based flow field with sparse flow values from $d$. Constraints are then interpolated into each voxel of the original unsmoothed segmentation map hull, providing the final constraints for integration into the 3D optical flow network.
@@
Our 3D optical flow network builds on the ARFlow [9] baseline, which employs a PWC-Net-based model [22] adapted to 3D imaging  [8]. The core objective of our 3D optical flow network is expressed in the following equation:
@@
The $\mathcal{L}_{1}$ loss captures the pixel-wise intensity disparities, ensuring accurate matching between the target and warped images. $\mathcal{L}_{ssim}$ [23] focuses on preserving structural details and perceptual quality during optimization.
@@ 4.2.1 Optimization
Our optimization scheme comprises pre-training on raw cardiac 4DCT scans, followed by optimization for a single pair of images from our training data, using $\mathcal{L}_{1}$ and $\mathcal{L}_{ssim}$. We then further optimize the network by incorporating $\mathcal{L}_{constraints}$.
@@ 5 Experiments and Results
Our experimental evaluations consisted of in-depth assessments of the efficacy of our method through optimization on a synthetic dataset enriched with ground truth annotations for LV 3D optical flow. The dataset consisted of 300 unique scan pairs, each exhibiting torsion ranging equally from 0 to approximately 35 degrees and distributed across various torsion centers as described in the supplementary material. The scans were taken from various sources [4][27][11][24][20].
We compared the performance of our method against two baseline approaches. The first baseline, here termed ”3D ARFlow”, corresponds to the network before optimization with additional constraint flow, akin to [8]. The second baseline, akin to DeepStrain [14], involved 3D ARFlow with an additional anatomical loss that encouraged maximal overlap between the predicted and ground truth segmentations. Both baselines were trained for 2500 iterations.
@@ 5.1 Results
(b) Angular
Error
(c) Locally-Radial
@@
(d) Locally-
Tangential mEPE
(e) Radial
@@
(f) Circumferential
mEPE
(g) Longitudinal
mEPE
Figure 3: Comprehensive comparison of proposed model and two baselines in accurately capturing cardiac deformation across a range of torsion angles. The metrics evaluated included the mEPE and mean angular error over the myocardium volume, the mEPE in locally-radial and locally-tangential directions within the segmentation hull, and the component-wise mEPE in radial, circumferential, and longitudinal directions within the myocardium volume. As depicted, our model exhibited strengths that were particularly evident in 3(a) representing the overall mEPE, and pronounced advantages in locally-tangential components 3(d), which aligns with our method’s goal of addressing the aperture problem, a challenge that was difficult to face by the two baseline models. Error bars represent the standard errors (SE).
Subfigure 3(a) shows that as the torsion angle increased, the mEPE across the myocardium region in our method consistently exhibited a reduction in comparison to the baselines.
Subfigure 3(b) depicts the quantification of the dissimilarity between the predicted and true flow directions based on the angular error. Our method outperformed the baselines in particular with increasing torsion angles, thus highlighting its effectiveness in capturing accurate directional information. In scenarios with lower torsion angles, our model experienced a relative decrease in performance in the angular error metric, possibly due to the increased sensitivity of angles in small ground truth displacement scenarios.
In terms of the locally-radial and locally-tangential flow components within the segmentation hull (Subfigures 3(c)3(d)), our method demonstrated superior performance as the torsion angle increased, particularly in capturing the intricacies of locally-tangential flow patterns, as compared to the baseline models. Our model exhibited significant advantages in locally-tangential components, and its performance in locally-radial aspects still remained competitive, with some variations observed in lower torsion ranges. This emphasizes the ability of our approach to effectively capture locally-tangential flow patterns within the anatomical regions of interest, by addressing the challenges associated with the aperture problem.
Breaking down the mEPE into radial, circumferential, and longitudinal components (Subfigures 3(e),3(f),3(g)) provides additional insights. Subfigures  3(e) and 3(f) consistently showcase our model’s superior performance across various torsion angles. Although there was a relative decrease in performance at lower torsion angles in the longitudinal component (Subfigure 3(g)), all three models, including ours, had substantially lower errors than for the other components.

### v2

@@ 1 Introduction
Traditional neural networks (NNs) have shown promise in analyzing cardiac movement from imaging data [4] through the use of designated architectures [14][16][20][26], optimization schemes [3], priors and constraints [27][6][2][15]. However, in addition to the scarcity of annotated high-resolution imagery, they also often suffer from the locality problem, since they struggle to predict tangential movements accurately (as depicted in figure 1(a)). While they excel at minimizing photometric energy to capture radial movements of the myocardium, their inherent locality limits their ability to comprehensively model subtle tangential movements. This limitation is referred to as the aperture problem, as illustrated in figure 1(b).
To address the challenge of capturing both radial and tangential components of myocardium motion, we propose a holistic approach that integrates DL techniques with spectral methods, and specifically FMs [17]. FMs, initially developed for non-rigid shape matching, operate by projecting functions from different shapes onto the eigenfunctions of the Laplacian Beltrami Operator (LBO) [22]. This LBO projection effectively captures the intrinsic geometry of the shapes, reducing the dimensionality of the problem and improving solution accuracy. By considering the LV surface as a manifold, we leverage FMs’ global perception for understanding myocardium motion.
Our comprehensive approach as detailed in this paper includes a 3D DL architecture and the incorporation of spectral methods. We leverage 2D constraints derived from spectral correspondence methods to significantly enhance the accuracy of surface mapping. These constraints provide a more robust foundation compared to traditional segmentation priors by precisely specifying the pixel correspondences between timesteps.
@@ 3 The Aperture Problem
Global approaches like Horn and Schunk [7] and Convolutional Neural Networks (CNNs) have been suggested but remain curtailed by the aperture problem. Despite CNNs’ broader perspective, their local context makes them prone to this issue. In cardiac CT imaging, this limitation is a major impediment because CNNs may fail to accurately estimate tangential motion on the dynamic LV myocardium surface due to their localized receptive fields.
@@ 4 Method
Figure 2: CardioSpectrum Architecture: The NN analyzes 3D image pairs of cardiac cycle timesteps, incorporating 2D constraints from ZoomOut. These constraints, derived from segmentations converted into meshes, result in a 3D optical flow.
@@ 4.1 Global Constraints from Functional Maps
To extract LV surface constraints, we use the ZoomOut method [13], which iteratively evaluates the FM matrix, transitioning between spatial and spectral domains. This process employs a progressive zooming-out strategy, gradually expanding the dimensionality of the functional mapping and refining point correspondences at each step, establishing correspondences between triangle mesh representations of two timesteps, derived from their smoothed LV segmentations.
We map these surface constraints to a 3D grid by calculating an array of 3D displacements, D, by subtracting corresponding vertex locations on the meshes. Using rounded vertex locations as indices, we fill a 3D grid-based flow field with sparse flow values from D. Constraints are then interpolated into each voxel of the original unsmoothed segmentation map hull, providing the final constraints for the 3D optical flow network.
@@
Our 3D optical flow network is built upon the ARFlow [10] training methodology and recommended model architecture, which employs a PWC-Net-based model  [23]. Adapted for 3D imaging as outlined in [9], our model integrates essential components including a feature pyramid extractor, cost volume, a flow estimator, and a refinement context network. The core objective of our 3D optical flow network is expressed in the following equation:
@@
The $\mathcal{L}_{1}$ loss captures the pixel-wise intensity disparities, ensuring accurate matching between the target and warped source images. $\mathcal{L}_{ssim}$ [24] focuses on preserving structural details and perceptual quality during optimization.
@@ 4.2.1 Optimization
Our optimization scheme involves two stages. First, we pre-train our model on a large motion-based dataset from raw cardiac 4DCT scans. Next, we fine-tune the model for each individual pair using $\mathcal{L}{1}$ and $\mathcal{L}{ssim}$, followed by further optimization with $\mathcal{L}_{constraints}$.
@@ 5 Experiments and Results
Our experimental evaluations consisted of in-depth assessments of the efficacy of our method through optimization on a synthetic dataset enriched with ground truth annotations for LV 3D optical flow. The dataset consisted of 300 unique scan pairs of systole and diastole, each exhibiting torsion ranging equally from 0 to approximately 35 degrees and distributed across various torsion centers. The data distribution and deformation process are described in [29]. The scans were taken from various sources [5][28][12][25][21].
We compared the performance of our method against two baseline approaches. The first baseline, here termed ”3D ARFlow”, corresponds to the network before optimization with additional constraint flow, akin to [9]. The second baseline, akin to [15][2], involved 3D ARFlow with an additional anatomical loss that encouraged maximal overlap between the predicted and ground truth segmentations. Both baselines were trained for 2500 iterations.
@@ 5.1 Results
(b) Locally-Radial
@@
(c) Locally-Tangential
@@
Figure 3: Comprehensive comparison of proposed model and two baselines in accurately capturing cardiac deformation across a range of torsion angles. The metrics evaluated included the mEPE over the myocardium volume, the mEPE in locally-radial and locally-tangential directions within the segmentation hull. As depicted, our model exhibited strengths that were particularly evident in 4(a) representing the overall mEPE, and pronounced advantages in locally-tangential components 4(c), which aligns with our method’s goal of addressing the aperture problem, a challenge that was difficult to face by the two baseline models. Error bars represent the standard errors (SE). See Fig. 1 in supplementary material for the component-wise mEPE in radial, circumferential, and longitudinal directions and mean angular error.
Subfigure 4(a) shows that as the torsion angle increased, the mEPE across the myocardium region in our method consistently exhibited a reduction in comparison to the baselines. In terms of the locally-radial and locally-tangential flow components within the segmentation hull (Subfigures 4(b)4(c)), our method demonstrated superior performance as the torsion angle increased, particularly in capturing the intricacies of locally-tangential flow patterns compared to the baseline models. Our model exhibited significant advantages in locally-tangential components, and its performance in locally-radial aspects still remained competitive, with some variations observed in lower torsion ranges. This emphasizes the ability of our approach to effectively capture locally-tangential flow patterns within the anatomical regions of interest, by addressing the challenges associated with the aperture problem. Breaking down the mEPE into radial, circumferential, and longitudinal components provides additional insights. For detailed analyses on these components, as well as the mean angular error, see the supplementary material.
@@ Acknowledgments.
Disclosure of Interests.
The authors have no competing interests to declare that are relevant to the content of this article.
Supplementary Material
(a) Radial mEPE
(b) Circumferential mEPE
(c) Longitudinal mEPE
(d) Angular Error
Figure 4: Additional comparison of the proposed model and two baselines in capturing cardiac deformation across torsion angles, breaking down mEPE into radial (4(a)), circumferential (4(b)), and longitudinal (4(c)) components. Radial and circumferential mEPE consistently showcase our model’s superior performance. Although performance relatively decreased at lower torsion angles in the longitudinal component, all models had lower errors than in other components.
Subfigure 4(d) shows dissimilarity between predicted and true flow directions based on angular error, with our method outperforming baselines particularly at higher torsion angles. Error bars represent standard errors (SE). These figures supplement Figure 3 and Section 5.1 of the main paper.
Table 1: Comparison of mEPE within the myocardium segmentation mask, along SE, at various torsion angles using CardioSpectrum and two baseline methods. Bold values denote significant improvement, and underlined values indicate comparable performance.
  Torsion [deg] mEPE $\pm$ SE [mm]
3D ARFlow 3D ARFlow + Anatomical Loss CardioSpectrum (Ours)
0 1.25 $\pm$ 0.043 0.99 $\pm$ 0.06 1.67 $\pm$ 0.13
2.9 1.44 $\pm$ 0.039 1.16 $\pm$ 0.04 1.73 $\pm$ 0.06
5.8 1.92 $\pm$ 0.11 1.60 $\pm$ 0.10 1.94 $\pm$ 0.08
8.7 2.67 $\pm$ 0.24 2.14 $\pm$ 0.17 2.26 $\pm$ 0.11
11.7 3.53 $\pm$ 0.39 2.80 $\pm$ 0.27 2.38 $\pm$ 0.10
14.6 4.56 $\pm$ 0.56 3.51 $\pm$ 0.38 2.59 $\pm$ 0.13
17.5 5.65 $\pm$ 0.71 4.52 $\pm$ 0.55 2.93 $\pm$ 0.15
20.4 6.90 $\pm$ 0.88 5.61 $\pm$ 0.72 3.08 $\pm$ 0.18
23.3 8.25 $\pm$ 1.08 6.93 $\pm$ 0.96 3.35 $\pm$ 0.19
26.2 9.72 $\pm$ 1.26 8.24 $\pm$ 1.14 3.59 $\pm$ 0.22
29.1 10.93 $\pm$ 1.43 9.51 $\pm$ 1.34 3.94 $\pm$ 0.27
32.1 12.26 $\pm$ 1.59 10.68 $\pm$ 1.50 4.29 $\pm$ 0.29
(a) Coordinate System for Cardiac Movement
(b) Visualization of Synthetic Deformation
(c) Segmentation Map Hull mEPE
Figure 5: (5(a)) The cardiac cycle involves radial (LV cavity center to myocardial wall), circumferential (tangential along the epicardial wall), and longitudinal (along the LV’s long axis) LV movements. Locally-tangential and locally-radial movements refer to projections onto and perpendicular to the myocardial surface’s tangent plane.
(5(b)) A sample before (top) and after (bottom) deformation. Arrows represent selected ground truth annotations. Sample from the 3D Slicer library (http://www.slicer.org).
(5(c)) Comparing mEPE of CardioSpectrum, baselines, and ZoomOut-derived constraints across torsion angles, evaluated over the segmentation map hull. ”Constraints” is composed of errors from ZoomOut and voxel-mesh conversions, impacting CardioSpectrum’s performance, especially at lower angles.